# Automatically Generating UI Code from Screenshot: A Divide-and-Conquer-Based Approach

### v2

@@ Abstract.
In this paper, we propose DCGen, a divide-and-conquer-based approach to automate the translation of webpage design to UI code. DCGen starts by dividing screenshots into manageable segments, generating descriptions for each segment, and then reassembling them into complete UI code for the entire screenshot. We conduct extensive testing with a dataset comprised of real-world websites and various MLLMs and demonstrate that DCGen achieves up to a 14% improvement in visual similarity over competing methods. Human evaluations show that DCGen can help developers implement webpages significantly faster and more similar to the UI designs. To the best of our knowledge, DCGen is the first segment-aware MLLM-based approach for generating UI code directly from screenshots.
@@
isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Developing a website’s graphical user interface (GUI) involves two successive activities: designing and implementing the user interface (UI). UI design contains graphic artists designing and customizing the website layout, which is then conveyed to developers in pixel-based concept drawings (Chen et al., 2018; Kuusinen and Mikkonen, 2013). Afterward, UI implementation focuses on converting the visual design into functional UI code using GUI frameworks.
@@
The challenges in GUI production necessitate an automated approach to translating UI designs (e.g., images, sketches) into GUI code, referred to as Design-to-Code. The automated design-to-code techniques offer benefits to individuals and companies. First, it has the potential to democratize front-end web application development (Nguyen and Csallner, 2015), allowing non-experts to easily build web applications. Second, it allows smaller companies to focus more on visual features, rather than on translating designs into functional code (Moran et al., 2018a).
@@
Despite its promising benefits, practical solutions for generating code implementation from website design have been underexplored. A line of research made efforts to transform GUI design into code for Android mobile apps but these approaches are not generalizable to the website design containing complicated interfaces (Chen et al., 2018; Nguyen and Csallner, 2015; Abdelhamid et al., 2020; Moran et al., 2018a). To enable finger-friendly touch targets, mobile apps utilize minimalist designs to avoid clutter on small screens, resulting in fewer visual elements and simpler structures. Additionally, mobile development relies on platform-specific tools (Jabangwe et al., 2018) (e.g., iOS uses Xcode, Android uses Android Studio), which do not apply to website development that typically uses HTML, CSS, and JavaScript (Sanghvi, 2024; kno, 2024; ink, 2024). Another line of research generates website UI code from screenshots using deep learning-based techniques (e.g., CNN, LSTM), but they rely on synthetic data. For example, Beltramelli et al. (Beltramelli, 2018) integrates CNN and RNN models to generate code from screenshots. However, this method is limited to a narrow range of simple user interface designs, accommodating only five types of web elements. Despite training on pairs of screenshots and code, DL-based approaches are still restricted by their training data which may contain limited web elements and programming knowledge. In a nutshell, existing studies are unsuitable for building real-world software websites.
@@
Multimodal Large Language Models (MLLMs) have seen significant advancements recently (Dai et al., 2023; OpenAI, [n. d.]; Google, 2024), offering an alternative solution for design-to-code. By integrating image processing capabilities into large language models (LLMs), MLLMs have demonstrated superior ability than CNNs for understanding images and answering visual questions (Yang et al., 2023). Furthermore, research has shown that LLMs have remarkable performance on various code intelligence tasks (Hou et al., 2023), including code generation (Yu et al., 2023; Li et al., 2023a; Du et al., 2023; Dong et al., 2023; Jiang et al., 2023; Gilbert et al., 2023), code completion (Nijkamp et al., 2022; Ding et al., 2023; Dibia et al., 2022; Li et al., 2023c, 2022; Chen et al., 2021), and code summarization (Mastropaolo et al., 2021, 2022; Gu et al., 2022; Chen et al., 2022a; Gao et al., 2023; Arakelyan et al., 2023). The strong image understanding and code generation capabilities of MLLMs showcase their potential to transform website screenshots into functional web code.
@@
To understand MLLMs’s capability in transforming screenshots to UI code, we conduct a motivating study to identify three main issues in this process: (1) element omission, where certain visual components are missing in the generated code, (2) element distortion, where elements are inaccurately reproduced in terms of shape, size, or color, and (3) element misarrangement, meaning the elements are not positioned or ordered correctly relative to their design layout. In contrast, when we cropped the inaccurately generated images into smaller pieces and asked MLLMs to re-generate UI code for each segment, we observed enhanced performance on these smaller, cropped screenshots. In short, the smaller and more focused the image, the better the resulting code quality.
Building on this insight, we propose a novel Divide-and-Conquer-based method to Generate website code from UI screenshots, namely DCGen. This end-to-end approach follows the principles of a typical divide-and-conquer algorithm: breaking down a complex problem into smaller parts, solving each part individually, and then combining the solutions to address the original problem. Specifically, DCGen consists of two phases: division and assembly. The first phase involves slicing the screenshots into smaller, more manageable, yet semantically meaningful pieces, and generating the HTML and CSS code for each distinctive segment. The division aligns with the real-world front-end development practices (Chopper, 2024) and is achieved via a novel image segmentation algorithm. DCGen then generates descriptions for each segment, followed by recursively reassembling these descriptions to reconstruct the entire website via UI code. This reassembly process is the reverse of the division procedure, where code from smaller, child segments is progressively integrated to build up their parent segments. This assembly continues until the full website structure is restored.
Our study curated a dataset comprising 1,000 top-ranked real-world websites. We evaluate the effectiveness of our methodology across several cutting-edge MLLMs, using 120 webpages sampled from the dataset. DCGen achieved up to 14% improvement in visual similarity between the original and generated websites compared to other design-to-code methods. We further demonstrate that DCGen is robust against various webpage complexities. Furthermore, DCGen is shown to generalize well across different MLLMs, demonstrating the effectiveness of the divide-and-conquer methodology. Human evaluations show that DCGen can produce more visually similar target webpages than competing methods and can help developers implement webpages faster and more similar to the UI design images.
@@
  • We propose DCGen with a divide-and-conquer approach. It starts with generating descriptions for individual image segments, and then merges these solutions to produce the complete UI code.
@@ 2. Background
Let $C_{0}$ be a file of HTML+CSS code of a webpage, $I_{0}$ be the screenshot of the webpage, and $M$ be an MLLM. The design-to-code task takes the image $I_{0}$ as input and outputs a file of generated HTML+CSS code $C_{g}=M(I_{0})$ that approximate $C_{0}$. Figure 4 illustrates an example of the input and output. The quality of $C_{g}$ is evaluated based on both functional similarity and visual similarity. Specifically, $C_{g}$ should be functionally similar to $C_{0}$, meaning it consists of a similar set of HTML elements and a comparable nested structure of the elements. Additionally, $I_{g}$, the screenshot of the webpage rendered from $C_{g}$, should be visually similar to $I_{0}$.
@@ 3. Dataset Collection
This section illustrates how we develop a dataset containing representative websites for this study. We randomly sample websites from the top 2,000 listed on the Tranco [^1] https://tranco-list.eu/. Accessed in Jan 2024  (Pochat et al., 2018), and process them until we have gathered 1,000 valid websites. First, we filter out all invalid websites (e.g., blocked, empty, or requiring human verification). Next, to make the websites self-contained and eliminate external dependencies that affect the UI code generation, we save each one into a single HTML file using the SingleFile toolkit [^2] https://github.com/gildas-lormeau/SingleFile, remove all external links, and replace images with placeholders. Finally, we clean the websites by removing all elements that do not impact the website’s appearance (i.e., hidden elements). The dataset collection yields 1,000 websites containing an average of 26 unique tags, where each HTML code file has an average length of 102,859 tokens.
Table 1 provides quantitative metrics to evaluate the complexity and comprehensiveness of our dataset. (1) Length: We tokenize the HTML code files in each category using the GPT-2 tokenizer. The average number of tokens is 102,859 (Min: 338, Max: 992,008, Std: 135,999), which suggests that our dataset covers a wide range of HTML code complexity. (2) Total number of tags: We counted the number of tags in HTML files, finding an average of 964 (Min: 1, Max: 16,331, Std: 1,370), which provides insights into the structural complexity of the documents. (3) DOM tree depth: We calculated the Document Object Model (DOM) tree depth in HTML files. The average depth is 17 (Min: 1, Max: 40, Std: 7), indicating the nesting complexity of HTML tags. (4) Number of unique tags: We also counted the number of unique tags, with an average of 26 (Min: 1, Max: 86, Std: 11), reflecting the diversity and content richness of the HTML files. Compared with existing datasets (Laurençon et al., 2024; Si et al., 2024; Beltramelli, 2018), our dataset is realistic and diverse in website complexity, making it well-suited for experiments in our paper.
@@
Length (tokens) 338 992,008 102,859
Tag Count 1 16,331 1,370
DOM Depth 1 40 17
Unique Tags 1 86 26
Total size 1,000
@@ 4.1. Study Preparation
This step identifies the bounding boxes for each element (tag) in the code within the webpage screenshot, referred to as visual elements. Using Selenium WebDriver [^3] https://selenium-python.readthedocs.io/, we extract the bounding box coordinates for all elements and eliminate any bounding boxes that intersect, as webpage elements typically do not overlap. This collection of bounding boxes enables us to locate all HTML elements in a website’s screenshot. Overall, these webpages contain a total of 1,699 HTML elements, with individual webpages featuring between 4 and 434 elements, averaging 142 elements per site.
Figure 1. Annotation tool. Annotators are provided with an original webpage (left) with a bounding box specifying the visual element and a generated webpage (right).
@@ 4.2. What failures does MLLM produce when converting design to code?
We recruit two graduate students in our university as annotators to assess the accuracy of visual elements in MLLM-generated webpages compared to the original ones. Figure 1 illustrates the annotation interface. The annotators are provided with pairs of screenshots, with the original on the left and the MLLM-generated version on the right. A red bounding box highlights the visual element in the original screenshot. Annotators are asked with two tasks: (1) Draw the bounding box around the corresponding element in the MLLM-generated screenshot. (2) Describe any differences between the generated visual element and the original in short phrases.
@@
Category Percentage (%) Count
Omission 85.34 1,450
Distortion 2.56 44
Misarrangement 12.71 216
Correct 2.35 40
Total 1,699
@@ 4.3. Can we correct mistakes by visual segments?
After noting the unsatisfactory performance in generating precise UI code, we became interested in understanding the reasons for these errors. LLMs are known to excel at simple, small-step tasks, and struggle with long-term planning for complex tasks (Wei et al., 2022b). To address this, existing studies have developed the Chain-of-Thought paradigm or LLM Agents, which decompose complex assignments into smaller, manageable sub-tasks, enabling models to perform more reasoning steps  (Feng and Chen, 2024).
@@
The first case focuses on omitted elements. We isolate each missing element by cropping the screenshot to include only that specific element and then request the model to generate the corresponding code. Figure 3 depicts this process, where Figure 3b is the Frequently Asked Questions (FAQ) area generated from the full screenshot and Figure 3c is the FAQ area generated from the cropped screenshot of the original FAQ area. The results show that with these isolated inputs, the model successfully generates the previously omitted content. This demonstrates that the model is intrinsically capable of producing the required outputs but struggles with complex inputs.
@@ 5. Methodology
In this section, we present DCGen, an end-to-end framework for automated UI code generation from screenshots. Motivated by the traditional “divide and conquer” algorithmic technique, DCGen decomposes a complicated screenshot into smaller, doable visual segments, solves each part individually, and then combines the solutions for the original problem.
DCGen accepts website screenshots and outputs a file of UI Code. Figure 4 shows the overall workflow of DCGen, consisting of two stages: division and assembly. In the division stage, DCGen segments the screenshot recursively into smaller pieces. Specifically, DCGen first searches for horizontal separation lines in a webpage and splits the image accordingly. For each horizontal segment, DCGen then looks for vertical separation lines and further divides the segment. This recursive “horizontal then vertical splitting” process continues, breaking the image into smaller, complete pieces until no more separation lines are found or a user-defined maximum depth is reached. At this point, DCGen passes each image segment to MLLMs for code generation. Afterward, in the assembly stage, code from smaller, child segments is progressively integrated to build up their parent segments. This recursive assembly continues until the full structure of the website is restored.
@@ 5.1. Division
The following section delineates a novel division algorithm, which comprises two main components: separation line detection and screenshot subdivision. The separation line detector first identifies all explicit or implicit separation lines within screenshots (either vertical or horizontal). Then, the screenshot subdivision component leverages these lines to segment the image into smaller and hierarchical segments. The two components work alternately, dividing the image horizontally, then vertically for the subimages, and again horizontally for the sub-subimages. Figure 3 illustrates various outcomes of the screenshot segmentation algorithm at different depths (i.e., the number of alternations).
@@ 5.1.1. Separation Line Detection
We observe that websites contain explicit (visible lines separating content) or implicit (blank spaces or borders between content elements) separation lines to organize different sections and contents. Figure 6 depicts examples of both types. The left figure employs explicit lines to separate the grey area from the purple area, while the right figure adapts implicit lines to arrange text sections.
@@
For detecting implicit separation lines, our approach incorporates a sliding window technique (lines 3-5), analyzing groups of consecutive rows to detect zones of minimal brightness variation, indicating potential blank spaces (lines 6-7). The algorithm compares each window against the rows directly above and below (lines 8-11). A separation line is identified if (1) the variance within the window is lower than a set threshold, suggesting a blank area, and (2) the brightness difference between this window and the rows directly above and below exceeds another threshold over more than 30% of the row’s length (line 12). This dual-threshold method facilitates the identification of subtle borders and blanks, recognizing both the low-content areas and the significant transitions at their edges. The thresholds and parameters used in the algorithm are manually optimized on a subset of our dataset.
For detecting explicit separation lines, the method is equivalent to finding implicit separation lines by setting the sliding window size to one. In particular, the algorithm scrutinizes each row for uniformity in pixel brightness. A row with variability below a predefined threshold is flagged as a potential line. Confirmation of a separation line occurs if the average difference in brightness between this row and its predecessor or successor surpasses a set threshold over at least 30% of the row.
@@
21:  return  sorted(lines)
(a) Original image
(b) Depth = 1
(c) Depth = 2
Figure 5. Examples of image subdivisions of different depths.
(d) Explicit.
(e) Implicit.

### v3

# Divide-and-Conquer: Generating UI Code from Screenshots

@@ Abstract.
In this paper, we propose DCGen, a divide-and-conquer-based approach to automate the translation of webpage design to UI code. DCGen starts by dividing screenshots into manageable segments, generating code for each segment, and then reassembling them into complete UI code for the entire screenshot. We conduct extensive testing with a dataset comprised of real-world websites and various MLLMs and demonstrate that DCGen achieves up to a 15% improvement in visual similarity and 8% in code similarity for large input images. Human evaluations show that DCGen can help developers implement webpages significantly faster and more similar to the UI designs. To the best of our knowledge, DCGen is the first segment-aware MLLM-based approach for generating UI code directly from screenshots.
@@
copyright: ccdoi: 10.1145/3729364journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE094publicationmonth: 7ccs: Software and its engineering Automatic programmingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Developing a website’s graphical user interface (GUI) involves two sequential stages: UI design and UI implementation. During the design stage, graphic artists create and customize the website layout, which is then provided to developers as pixel-based concept drawings (Chen et al., 2018; Kuusinen and Mikkonen, 2013). Subsequently, the implementation stage involves translating these visual designs into functional UI code using GUI frameworks. Screenshots of the GUI can effectively approximate the original UI designs, as they capture the intended pixel-level layout and visual structure of the design before being converted into operational code.
@@
The challenges in GUI production necessitate an automated approach to translating UI designs (e.g., images, Photoshop documents) into GUI code, referred to as Design-to-Code. The automated design-to-code techniques offer benefits to individuals and companies. First, it has the potential to democratize front-end web application development (Nguyen and Csallner, 2015), allowing non-experts to easily build web applications. Second, it allows smaller companies to focus more on visual features, rather than on translating designs into functional code (Moran et al., 2018a).
@@
Despite its promising benefits, practical solutions for generating code implementation from website design have been underexplored. A line of research made efforts to transform GUI design into code for Android mobile apps but these approaches are not generalizable to the website design containing complicated interfaces (Chen et al., 2018; Nguyen and Csallner, 2015; Abdelhamid et al., 2020; Moran et al., 2018a). To enable finger-friendly touch targets, mobile apps utilize minimalist designs to avoid clutter on small screens, resulting in fewer visual elements and simpler structures. Additionally, mobile development relies on platform-specific tools (Jabangwe et al., 2018) (e.g., iOS uses Xcode, Android uses Android Studio), which do not apply to website development that typically uses HTML, CSS, and JavaScript (Sanghvi, 2024; kno, 2024; ink, 2024). Another line of research generates website UI code from screenshots using deep learning-based techniques (e.g., CNN, LSTM), but they rely on synthetic data. For example, Beltramelli et al. (Beltramelli, 2018) integrates CNN and RNN models to generate code from screenshots. However, this method is limited to a narrow range of simple user interface designs, accommodating only five types of web elements. Despite training on pairs of screenshots and code, DL-based approaches are still restricted by their training data, which may contain limited web elements and programming knowledge. In a nutshell, existing studies are unsuitable for building real-world software websites.
@@
Multimodal Large Language Models (MLLMs) have seen significant advancements recently (Dai et al., 2023; OpenAI, [n. d.]; Google, 2024), offering an alternative solution for design-to-code. By integrating image processing capabilities into large language models (LLMs), MLLMs have demonstrated superior ability than CNNs for understanding images and answering visual questions (Yang et al., 2023). Furthermore, research has shown that LLMs have remarkable performance on various code intelligence tasks (Hou et al., 2023), including code generation (Yu et al., 2023; Li et al., 2023a; Du et al., 2023; Dong et al., 2023; Jiang et al., 2023; Gilbert et al., 2023), code completion (Nijkamp et al., 2022; Ding et al., 2023; Dibia et al., 2022; Li et al., 2023c, 2022; Chen et al., 2021), and code summarization (Mastropaolo et al., 2021, 2022; Gu et al., 2022; Chen et al., 2022a; Gao et al., 2023; Arakelyan et al., 2023). The strong image understanding and code generation capabilities of MLLMs showcase their potential to transform website design images into functional web code.
@@
To understand MLLMs’s capability in transforming UI images to UI code, we conduct a motivating study to identify three main issues in this process: (1) element omission, where certain visual components are missing in the generated code, (2) element distortion, where elements are inaccurately reproduced in terms of shape, size, or color, and (3) element misarrangement, meaning the elements are not positioned or ordered correctly relative to their design layout. In contrast, when we cropped the inaccurately generated images into smaller pieces and asked MLLMs to re-generate UI code for each segment, we observed enhanced performance on these smaller, cropped UI images. In short, the smaller and more focused the image, the better the resulting code quality.
Building on this insight, we propose a novel Divide-and-Conquer-based method to Generate website code from UI designs, namely DCGen. This end-to-end approach follows the principles of a typical divide-and-conquer algorithm: breaking down a complex problem into smaller parts, solving each part individually, and then combining the solutions to address the original problem. Specifically, DCGen consists of two phases: division and assembly. The first phase involves slicing the design images into smaller, more manageable, yet semantically meaningful pieces, and generating the HTML and CSS code for each distinctive segment. The division aligns with the real-world front-end development practices (Chopper, 2024) and is achieved via a novel image segmentation algorithm. DCGen then generates code for each segment, followed by recursively reassembling the code to reconstruct the entire website. This reassembly process is the reverse of the division procedure, where code from smaller, child segments is progressively integrated to build up their parent segments. This assembly continues until the full website structure is restored.
Our study curated a dataset comprising 348 top-ranked real-world websites. We evaluate the effectiveness of our methodology across several cutting-edge MLLMs, using 111 webpages sampled from the dataset. DCGen achieved up to 8% improvement in visual similarity and 15% improvement in code similarity for large input images compared to other design-to-code methods. We further demonstrate that DCGen is robust against complex webpages. Moreover, DCGen is shown to generalize well across different MLLMs, demonstrating the effectiveness of the divide-and-conquer methodology. Human evaluations show that DCGen can produce more visually similar target webpages than competing methods and can help developers implement webpages faster.
@@
  • We propose DCGen with a divide-and-conquer approach. It starts with generating code for individual image segments, and then merges these solutions to produce the complete UI code.
@@ 2. Background
This work addresses the design-to-code task, where the input is a visual design of a webpage, and the goal is to generate HTML+CSS code that accurately reproduces it. Let $I_{0}$ be the design image of a webpage, and $C_{0}$ be the ideal HTML+CSS code. Given $I_{0}$, an MLLM $M$ generates code $C_{g}=M(I_{0})$, and the rendered output $I_{g}$ of $C_{g}$ should closely match $I_{0}$ in both structure and appearance.
Because design files are typically unavailable, we use screenshots of existing webpages as proxies for design images $I_{0}$, and use the corresponding HTML+CSS code as ground truth $C_{0}$. This allows us to construct datasets for evaluating the model’s ability to recover code from visual designs. The quality of $C_{g}$ is assessed based on (1) code similarity to $C_{0}$—matching elements and structure—and (2) visual similarity between the rendered output $I_{g}$ and the design image $I_{0}$.
@@ 3. Dataset Collection
This section illustrates how we develop a dataset containing representative websites for this study. We randomly sample websites from the top 500 listed on the Tranco [^1] https://tranco-list.eu/. Accessed in Jan 2024  (Pochat et al., 2018), and process them. First, we filter out all invalid websites (e.g., blocked, empty, or requiring human verification). Next, to make the websites self-contained and eliminate external dependencies that affect the UI code generation, we save each one into a single HTML file using the SingleFile toolkit [^2] https://github.com/gildas-lormeau/SingleFile, remove all external links, and replace images with placeholders. Finally, we clean the websites by removing all elements that do not impact the website’s appearance (i.e., hidden elements). The dataset collection yields 348 websites containing an average of 26 unique tags, where each HTML code file has an average length of 99,431 tokens.
Table 2 provides quantitative metrics to evaluate the complexity and comprehensiveness of our dataset. (1) Length: We tokenize the HTML code files in each category using the GPT-2 tokenizer. The average number of tokens is 99,431, which suggests that our dataset covers a wide range of HTML code complexity. (2) Total number of tags: We counted the number of tags in HTML files, finding an average of 961, which provides insights into the structural complexity of the documents. (3) DOM tree depth: We calculated the Document Object Model (DOM) tree depth in HTML files. The average depth is 17, indicating a high nesting complexity of HTML tags. (4) Number of unique tags: We also counted the number of unique tags, with an average of 26, reflecting the diversity and content richness of the HTML files.
@@
Length (tokens) 383 868,108 99,431
Tag Count 2 16,331 961
DOM Depth 2 40 17
Unique Tags 2 66 26
Total size 348
Category Percentage (%) Count
Omission 85.34 1,450
Distortion 2.56 44
Misarrangement 12.71 216
Correct 2.35 40
Total 1,699
@@ 4.1. Study Preparation
This step identifies the bounding boxes for each element (tag) in the code within the webpage screenshot, referred to as visual elements. Using Selenium WebDriver [^3] https://selenium-python.readthedocs.io/, we extract the bounding box coordinates for all elements and eliminate any bounding boxes that intersect. This collection of bounding boxes enables us to locate all HTML elements in a website’s screenshot. Overall, these webpages contain a total of 1,699 HTML elements, with individual webpages featuring between 4 and 434 elements, averaging 142 elements per site.
@@ 4.2. What failures does MLLM produce when converting design to code?
We recruit two graduate students in our university as annotators to assess the accuracy of visual elements in MLLM-generated webpages compared to the original ones. The annotators are provided with pairs of screenshots, with the original on the left and the MLLM-generated version on the right. A red bounding box highlights the visual element in the original screenshot. Annotators are asked with two tasks: (1) Draw the bounding box around the corresponding element in the MLLM-generated screenshot. (2) Describe any differences between the generated visual element and the original in short phrases.
@@ 4.3. Can we correct mistakes by visual segments?
After noting the unsatisfactory performance in generating precise UI code, we investigate the reasons for these errors. LLMs are known to excel at simple, small-step tasks, and struggle with long-term planning for complex tasks (Wei et al., 2022b). To address this, existing studies have developed the Chain-of-Thought paradigm or LLM Agents, which decompose complex assignments into smaller, manageable sub-tasks, enabling models to perform more reasoning steps  (Feng and Chen, 2024).
@@
The first case focuses on omitted elements. We isolate each missing element by cropping the screenshot to include only that specific element and then request the model to generate the corresponding code. Figure 2 depicts this process, where Figure 2b is the Frequently Asked Questions (FAQ) area generated from the full screenshot, and Figure 2c is the FAQ area generated from the cropped screenshot. The results show that with these isolated inputs, the model successfully generates the previously omitted content. This demonstrates that the model is intrinsically capable of producing the required outputs but struggles with complex inputs.
@@ 5. Methodology
In this section, we present DCGen, an end-to-end framework for automated UI code generation from design images. Motivated by the traditional “divide and conquer” algorithmic technique, DCGen decomposes a complicated screenshot into smaller, doable visual segments, solves each part individually, and then combines the solutions for the original problem.
DCGen accepts website screenshots and outputs a file of UI Code. Figure 3 shows the overall workflow of DCGen, consisting of two stages: division and assembly. In the division stage, DCGen searches for horizontal and vertical separation lines in an interleaved fashion and splits the image accordingly, breaking the image into smaller pieces until no more separation lines are found or a user-defined maximum depth is reached. At this point, DCGen passes each image segment to MLLMs for code generation. Afterward, in the assembly stage, code from smaller, child segments is progressively integrated to build up their parent segments. This recursive assembly continues until the full structure of the website is restored.
@@ 5.1. Division
The following section delineates a novel division algorithm, which comprises two main components: separation line detection and screenshot subdivision. The separation line detector first identifies all explicit or implicit separation lines within screenshots (either vertical or horizontal). Then, the screenshot subdivision component leverages these lines to segment the image into smaller and hierarchical segments. The two components work alternately, dividing the image horizontally, then vertically for the subimages.
@@ 5.1.1. Separation Line Detection
We observe that websites contain explicit (visible lines separating content) or implicit (blank spaces or borders between content elements) separation lines to organize different sections and contents. Figure 4 depicts examples of both types. Figure 4a employs explicit lines to separate the grey area from the purple area, while Figure 4b adapts implicit lines to arrange text sections.
@@
For detecting implicit separation lines, our approach incorporates a sliding window technique (lines 3-5), analyzing groups of consecutive rows to detect zones of minimal brightness variation, indicating potential blank spaces (lines 6-7). The algorithm compares each window against the rows directly above and below (lines 8-11). A separation line is identified if (1) the variance within the window is lower than a set threshold, suggesting a blank area, and (2) the brightness difference between this window and the rows directly above and below exceeds another threshold over more than portion_thr of the row’s length (line 9, 11). This dual-threshold method facilitates the identification of subtle borders and blanks, recognizing both the low-content areas and the significant transitions at their edges. The selection of thresholds and parameters are detailed in Section 7.1.
For detecting explicit separation lines, the method is equivalent to finding implicit separation lines by setting the sliding window size to one. In particular, the algorithm scrutinizes each row for uniformity in pixel brightness. A row with variability below a predefined threshold is flagged as a potential line. Confirmation of a separation line occurs if the average difference in brightness between this row and its predecessor or successor surpasses a set threshold over at least portion_thr of the row.
@@
21:
22:return  sorted(lines)
(a) Explicit.
(b) Implicit.